# Measuring entanglement without local addressing in quantum many-body simulators via spiral quantum state tomography

### v2

@@ Abstract
Introduction
Accurate and efficient determination of quantum states is essential for advancing quantum technologies, including digital quantum computers Lloyd1996-br ; Preskill2018-yj ; Bharti2022-vb and analog quantum simulators Georgescu2014-im ; Altman2021-te . Historically, experimental physics research involved testing theoretical models by comparing their predictions with measured values of specific material properties, such as electrical conductivity, magnetic susceptibility, thermal characteristics, and more. In recent years, the evolution of quantum technologies has shifted the focus toward extracting more fundamental information about quantum states and exploring their inherent quantum entanglement. However, conventional methods for estimating quantum states, typically represented by wavefunctions or more generally by density matrices, require computational resources that grow exponentially with system size Paris2004-xq ; Hawkes2014-lu . This challenge has become increasingly urgent as a variety of quantum platforms—such as superconducting circuits Devoret2013-bg ; Wendin2017-io , photonic qubits Kok2007-ww , trapped ions Blatt2012-mn ; Haffner2008-tq , Rydberg atom arrays Adams2020-ut ; Browaeys2020-un , and cold atoms in optical lattices Gross2017-am ; Schafer2020-hn —have scaled up to unprecedented levels.
Given a density matrix $\rho$ representing a quantum state, the expectation value of any observable $M$ can be computed by taking the trace of the product of $\rho$ and $M$: $\langle M\rangle={\rm Tr}(\rho M)$. Conversely, the density matrix $\rho$ can be experimentally reconstructed from a set of measurements on various observables that, when taken together, provide sufficient information to uniquely determine the quantum state. This process is generally known as quantum state tomography (QST) Paris2004-xq ; Hawkes2014-lu . Given that $\rho$ is a $d\times d$ Hermitian matrix with trace 1, it has $d^{2}-1$ independent real parameters. Therefore, to uniquely reconstruct $\rho$, in principle one needs a set of $d^{2}-1$ expectation values of independent observables: $\langle M_{n}\rangle$ with $n=1,2,\dots,d^{2}-1$.
For $N$-qubit systems, where $d=2^{N}$, the Pauli matrices ($X,Y,Z$) along with the identity $I$ are typically used as a local (single-qubit) basis of Hermitian operators. A “tomographically complete” set of observables can then be constructed as the set of $14^{N}-$ 1 operators
@@ Introduction
where $i_{1}i_{2}\dots i_{N}$ is the base-4 representation of $n=1,2,\dots,4^{N}-1$ and $(\sigma^{0},\sigma^{1},\sigma^{2},\sigma^{3})=(I,X,Y,Z)$. Thus, the procedure of full tomography requires $3^{N}$ orthogonal measurement settings, each producing outcomes for $2^{N}$ different observables. It then involves averaging the outcomes over many repetitions of the measurements for each setting, upon preparing the target state each time, to estimate accurate expectation values. The exponential growth of the measurement cost limits full QST to small systems of several qubits; for example, more than half a million measurements were required to reconstruct an eight-particle entangled state with trapped ions, achieving a fidelity of $F\approx 0.72$ Haffner2005-zv . To mitigate the practical difficulties, various techniques, each specialized in a different way, have been proposed for specific target states, e.g those states that are permutationally invariant Toth2010-jx ; Schwemmer2014-sp , represented by low-rank density matrices Gross2010-mz ; gross-11 ; NIPS2011_e820a45f ; Kalev2015-ur ; Schwemmer2014-sp , described by matrix product states Cramer2010-dx ; Lanyon2017-ab , or uniquely determined by their reduced density matrices Xin2017-xw .
@@
Particularly in the context of quantum simulations of condensed-matter physics and field theory, there is a significant demand for techniques that can reconstruct density matrices that have low rank, but represent a large number of qubits or particles, to study the many-body ground state of certain physically important Hamiltonians. The concept of compressed sensing for recovering a low-rank density matrix from an incomplete set of measurements offers an approach to significantly reduce the number of expectation values required for QST down to $\mathcal{O}(rd\log^{2}d)$, where $r$ denotes the rank of $\rho$ Gross2010-mz . However, in addition to the issue of the number of measurements, practical difficulties still arise in preparing an enormous number of measurement settings as the system size grows. In a standard implementation of QST, typically based on the Pauli basis (1), one usually needs to rotate local quantization axes into the $X$, $Y$, or $Z$ directions and then read out the qubits for each measurement setting. This requires precise one-by-one operations and local implementations, which become more complex and costly as the system scales up and integrates more densely.
@@
In this work, we propose an efficient QST protocol based on compressed sensing utilizing a “spiral” measurement set. A spiral measurement can be achieved by rotating qubits in a manner resembling a spin-spiral configuration observed in chiral magnets Kimura2007-gh ; Rybakov2016-qw , with a pitch angle $q$. Since this can be implemented using only the combination of global qubit rotations and the application of a longitudinal ($Z$) field gradient, it drastically reduces the cost of preparing the measurement settings required for QST and is especially beneficial for large-scale quantum systems that may not be completely suited for easy and clean single-qubit (-site) addressing, such as systems of ultra-cold atoms in optical lattices Gross2017-am ; Schafer2020-hn .
@@
Results
Quantum state tomography via spiral compressed sensing
Figure 1: Spiral measurements. a Illustrations of the three spiral measurement sets. The blue arrows indicate the local measurement axis at each qubit. b Pitch angle $q$ in the spiral plane of $\tilde{M}^{XY}(q)$.
We first introduce the protocol for QST via compressed sensing using a set of spiral measurement operators (Fig. 1a). Conventional compressed-sensing QST Gross2010-mz ; gross-11 ; NIPS2011_e820a45f ; Kalev2015-ur reconstructs a pure or nearly-pure quantum state, represented by a low-rank density matrix, using $\mathcal{O}(rd\log^{2}d)$ random samples from the Pauli measurements (1). The idea of the spiral compressed sensing is to use measurements based on three sets of spiral-spin operators, instead of Pauli operators:
@@ Results
where $N$ qubits (or sites) are supposed to be arranged along the $x$-axis at equal intervals, indexed by $i=1,2,\dots,N$. Here, $q$ and $i_{0}$ represent the pitch and origin of the spiral structure of the measurement axis (see Fig. 1b). Measuring a spiral operator, say $\tilde{M}^{XY}(q)$, on a given quantum state, or equivalently measuring $Z^{(i)\prime}$ on each qubit, can be carried out as follows: First, $Z$-axis rotation gates $R_{Z}(\theta_{i})$ are applied to every qubit with linearly varying angles with pitch $q$: $\theta_{i}=-q(i-i_{0})$. Subsequently, the $X$-basis measurements are performed for all the qubits by applying the Hadamard gates to them, followed by measurements in the computational basis ($Z$-basis) in a standard way. The other spirals, $\tilde{M}^{YZ}(q)$ and $\tilde{M}^{ZX}(q)$, can also be measured similarly by applying additional uniform gates to change the spiral plane into the $xy$-plane before the $R_{Z}(\theta_{i})$ rotations.
@@
where $O_{i}$ denotes the single-qubit operator $O$ acting on site $i$, for a short time interval of $\Delta t=q/B$. The spiral measurement setups with three different planes (shown in Fig. 1a) can then be prepared by combining this with globally uniform $\pi$/2 Rabi pulses Fukuhara2015-sn ; Mazurenko2017-yb ; Brown2017-fp . This process can also be viewed as the inverse of creating a spin-spiral state, starting from the initial “all-up” state Koschorreck2013-ur ; Bardon2014-bs ; Hild2014-mz ; Jepsen2020-ze ; Jepsen2021-xl .
By tuning the pitch $q$ through the manipulation of the strength $B$ and/or the application time $\Delta t$ of the field gradient, different measurement setups can be prepared for each spiral plane. When denoting the number of different pitch angles by $n_{q}$, the number of independent experimental setups is $3n_{q}$, where the factor of 3 comes from the three spiral planes. For each setup, $12^{N}-$ 1 mutually independent expectation values can be obtained, as it is possible to choose whether to include the measurement of each individual qubit in the calculation of the multi-site correlation functions. In other words, from a single setup, say $\tilde{M}^{XY}(q)$ with a specific $q$, one can simultaneously obtain $\langle\tilde{M}^{XY}(q)\rangle$ and all the correlation functions where any of its $Z^{\prime}_{i}$s are replaced with $I$. We denote the corresponding operators by $\tilde{M}^{\alpha\beta}_{\bm{a}}(q)$, where $\alpha\beta=XY,YZ,ZX$ and $\bm{a}\in\{0,1\}^{N}$ is a multi-index indicating the positions of the identity matrices (labeled by 0), e.g., $\tilde{M}^{XY}_{(1,0,\ldots,0)}(q)=Z^{(1)\prime}\otimes I\cdots\otimes I/\sqrt{d}$, $\tilde{M}^{YZ}_{(1,1,0,\ldots,0)}(q)=Z^{(1)\prime\prime}\otimes Z^{(2)\prime\prime}\otimes I\cdots\otimes I/\sqrt{d}$, $\tilde{M}^{ZX}_{(1,\ldots,1)}(q)=Z^{(1)\prime\prime\prime}\otimes\cdots\otimes Z^{(N)\prime\prime\prime}/\sqrt{d}=\tilde{M}^{ZX}(q)$, etc. Note that this family includes the product of only identities at all sites, which contains no information. Consequently, the number of measurements that can be obtained from the three spiral sets with $n_{q}$ different pitch angles is, in total, $\sim 3n_{q}2^{N}$, which can be made comparable to $\mathcal{O}(rN^{2}2^{N})$, the required number of Pauli measurements for successful compressed sensing Gross2010-mz , by choosing a sufficiently large $n_{q}$.
@@
Numerical tests
To demonstrate the efficiency and accuracy of the spiral QST, we test the reconstruction of random pure states ($r=1$) and rank-3 mixed states ($r=3$). For the pitch angle $q$, we employ $n_{q}=2N$ values with intervals of $\pi/N$, defined as
@@
The above parameters, while reasonable from a physical standpoint, are by no means the only possible choice. We utilize the singular value thresholding (SVT) algorithm Gross2010-mz ; Cai2008-zn for compressed sensing, sampling expectation values of spiral operators in the set $\{\tilde{M}^{\alpha\beta}_{\bm{a}}(q)\}$ associated to randomly chosen pitch angles from Eq. (4) (see Methods). Target states $\rho$ are generated by randomly selecting from the Haar measure on a $d\times r$-dimensional system ($d=2^{N}$) and tracing out the $r$-dimensional ancilla. To model typical noise during the preparation of the target state, we add depolarizing noise of strength $\gamma$ after generating the target states as $\rho\rightarrow(1-\gamma)\rho+\gamma I^{\otimes N}/d$. We also account for measurement noise modeled by Gaussian fluctuations with standard deviation $\sigma$ in calculating expectation values for each measurement: $\langle\tilde{M}^{\alpha\beta}_{\bm{a}}(q)\rangle\rightarrow\mathcal{N}(\mu=\langle\tilde{M}^{\alpha\beta}_{\bm{a}}(q)\rangle,\sigma)$
Figure 2: Numerical tests of spiral QST. Sample-averaged fidelity and trace distance for compressed-sensing tomography of (a) random pure states (rank $r=1$) and (b) random mixed states (rank $r=3$) of 8 qubits, plotted as functions of the number of measurements $m$ scaled by $d^{2}=4^{8}$. The sampled states include 5 percent depolarizing noise ($\gamma=0.05$) and measurement Gaussian statistical noise with a standard deviation of $\sigma=0.1/d$. Spiral and Pauli measurements are compared in each case. For $r=3$, the results of Pauli-based tomography were also reported in previous work Gross2010-mz .
@@
$T(\rho_{1},\rho_{2})=||\rho_{1}-\rho_{2}||_{1}/2,$
where $||A||_{1}\equiv{\rm Tr}(\sqrt{A^{\dagger}A})$ is the trace norm (or nuclear norm). These metrics take values 1 and 0, respectively, if and only if the two states $\rho_{1}$ and $\rho_{2}$ are identical. In Figs. 2a and 2b, we present $F(\rho,\rho^{\rm rec})$ and $T(\rho,\rho^{\rm rec})$ between the target state $\rho$ and the reconstructed state $\rho^{\rm rec}$ for pure and mixed states, respectively, of $N=8$ qubits, plotted as functions of the number of spiral measurements (expectation values), $m$, used in the SVT algorithm, expressed in units of $d^{2}$. The results include noise levels corresponding to $\gamma=0.05$ and $\sigma=0.1/d$, matching those used in the demonstrations with Pauli-basis measurements in Ref. Gross2010-mz .
As seen from the results for $r=1$, the performance of spiral QST is comparable to (though slightly worse than) the conventional Pauli-basis compressed sensing Gross2010-mz . The fidelity and trace distance indicate that the state reconstruction is very successful with only about 10 percent of a tomographically complete set of measurements. For mixed states with rank $r=3$, the results are also notable. Given that spiral measurements do not require precise single-site addressing nor a large number of experimental setups, this outcome highlights the practicality of our approach. Building on these results for random quantum states, we now aim to explore more realistic scenarios reflecting actual experimental conditions in the remainder of the paper.
Simulation of typical experimental situations

### v3

@@ Abstract
I Introduction
Accurate and efficient determination of quantum states is essential for advancing quantum technologies, including digital quantum computers [1, 2, 3] and analog quantum simulators [4, 5]. Historically, experimental physics research involved testing theoretical models by comparing their predictions with measured values of specific material properties, such as electrical conductivity, magnetic susceptibility, thermal characteristics, and more. In recent years, the evolution of quantum technologies has shifted the focus toward extracting more fundamental information about quantum states and exploring their inherent quantum entanglement. However, conventional methods for estimating quantum states, typically represented by wavefunctions or more generally by density matrices, require computational resources that grow exponentially with system size [6, 7]. This challenge has become increasingly urgent as a variety of quantum platforms—such as superconducting circuits [8, 9], photonic qubits [10], trapped ions [11, 12], Rydberg atom arrays [13, 14], and cold atoms in optical lattices [15, 16]—have scaled up to unprecedented levels.
Given a density matrix $\rho$ representing a quantum state, the expectation value of any observable $M$ can be computed by taking the trace of the product of $\rho$ and $M$: $\langle M\rangle={\rm Tr}(\rho M)$. Conversely, the density matrix $\rho$ can be experimentally reconstructed from a set of measurements on various observables that, when taken together, provide sufficient information to uniquely determine the quantum state. This process is generally known as quantum state tomography (QST) [6, 7]. Given that $\rho$ is a $d\times d$ Hermitian matrix with trace 1, it has $d^{2}-1$ independent real parameters. Therefore, to uniquely reconstruct $\rho$, in principle one needs a set of $d^{2}-1$ expectation values of independent observables: $\langle M_{n}\rangle$ with $n=1,2,\dots,d^{2}-1$.
For $N$-qubit systems, where $d=2^{N}$, the Pauli matrices ($X,Y,Z$) along with the identity $I$ are typically used as a local (single-qubit) basis of Hermitian operators. A “tomographically complete” set of observables can then be constructed as the set of $4^{N}-1$ operators
@@ I Introduction
where $i_{1}i_{2}\dots i_{N}$ is the base-4 representation of $n=1,2,\dots,4^{N}-1$ and $(\sigma^{0},\sigma^{1},\sigma^{2},\sigma^{3})=(I,X,Y,Z)$. Thus, the procedure of full tomography requires $3^{N}$ orthogonal measurement settings, each producing outcomes for $2^{N}$ different observables. It then involves averaging the outcomes over many repetitions of the measurements for each setting, upon preparing the target state each time, to estimate accurate expectation values. The exponential growth of the measurement cost limits full QST to small systems of several qubits; for example, more than half a million measurements were required to reconstruct an eight-particle entangled state with trapped ions, achieving a fidelity of $F\approx 0.72$ [17]. To mitigate the practical difficulties, various techniques, each specialized in a different way, have been proposed for specific target states, e.g those states that are permutationally invariant [18, 19], represented by low-rank density matrices [20, 21, 22, 23, 19], described by matrix product states [24, 25], or uniquely determined by their reduced density matrices [26]. It is worth mentioning a novel approach, namely shadow tomography [27, 28], which prescribes how to extract certain quantities of interest from a quantum state without reconstructing the whole density matrix, using relatively few Pauli measurements; available simulations show great potential, especially for the estimation of linear functions of the density matrix (expectation values of observables) [28].
@@
Particularly in the context of quantum simulations of condensed-matter physics and field theory, there is a significant demand for techniques that can reconstruct density matrices that have low rank, but represent a large number of qubits or particles, to study the many-body ground state of certain physically important Hamiltonians. The concept of compressed sensing for recovering a low-rank density matrix from an incomplete set of measurements offers an approach to significantly reduce the number of expectation values required for QST down to $\mathcal{O}(rd\log^{2}d)$, where $r$ denotes the rank of $\rho$ [20]. However, in addition to the issue of the number of measurements, practical difficulties still arise in preparing an enormous number of measurement settings as the system size grows. In a standard implementation of QST, typically based on the Pauli basis (1), one usually needs to rotate local quantization axes into the $X$, $Y$, or $Z$ directions and then read out the qubits for each measurement setting [29, 30]. This requires precise one-by-one operations and local implementations, which become more complex and costly as the system scales up and integrates more densely.
@@
In this work, we propose an efficient QST protocol based on compressed sensing utilizing a “spiral” measurement set. A spiral measurement can be achieved by rotating qubits in a manner resembling a spin-spiral configuration observed in chiral magnets [31, 32], with a pitch angle $q$. Since this can be implemented using only the combination of global qubit rotations and the application of a longitudinal ($Z$) field gradient, it drastically reduces the cost of preparing the measurement settings required for QST and is especially beneficial for large-scale quantum systems that may not be completely suited for easy and clean single-qubit (-site) addressing, such as systems of ultra-cold atoms in optical lattices [15, 16]. In fact, in the latter platform the single-site control of atoms is not impossible, but still challenging (especially when compared with trapped ions and atoms in optical tweezers), the fundamental reason being that the distance between qubits (spins) is similar to the size of “diffraction-limit” beams (a possible exception is given by spin-flip operations [33]).
@@
II Quantum state tomography via spiral compressed sensing
II.1 Methodology
Figure 1: Spiral measurements. (a) Illustrations of the three spiral measurement sets. The blue arrows indicate the local measurement axis at each qubit. (b) Pitch angle $q$ in the spiral plane of $\tilde{M}^{XY}(q)$.
We first introduce the protocol for QST via compressed sensing using a set of spiral measurement operators (Fig. 1(a)). Conventional compressed-sensing QST [20, 21, 22, 23] reconstructs a pure or nearly-pure quantum state, represented by a low-rank density matrix, using $\mathcal{O}(rd\log^{2}d)$ random samples from the Pauli measurements (1). The idea of the spiral compressed sensing is to use measurements based on three sets of spiral-spin operators, instead of Pauli operators:
@@ II.1 Methodology
where $N$ qubits (or sites) are supposed to be arranged along the $x$-axis at equal intervals, indexed by $i=1,2,\dots,N$. Here, $q$ and $i_{0}$ represent the pitch and origin of the spiral structure of the measurement axis [see Fig. 1(b)]. Measuring a spiral operator, say $\tilde{M}^{XY}(q)$, on a given quantum state, or equivalently measuring $Z^{(i)\prime}$ on each qubit, can be carried out as follows: First, $Z$-axis rotation gates $R_{Z}(\theta_{i})$ are applied to every qubit with linearly varying angles with pitch $q$: $\theta_{i}=-q(i-i_{0})$. Subsequently, the $X$-basis measurements are performed for all the qubits by applying the Hadamard gates to them, followed by measurements in the computational basis ($Z$-basis) in a standard way. The other spirals, $\tilde{M}^{YZ}(q)$ and $\tilde{M}^{ZX}(q)$, can also be measured similarly by applying additional uniform gates to change the spiral plane into the $xy$-plane before the $R_{Z}(\theta_{i})$ rotations.
@@
where $O_{i}$ denotes the single-qubit operator $O$ acting on site $i$, for a short time interval of $\Delta t=q/B$. The spiral measurement setups with three different planes [shown in Fig. 1(a)] can then be prepared by combining this with globally uniform $\pi$/2 Rabi pulses [34, 35, 36]. This process can also be viewed as the inverse of creating a spin-spiral state, starting from the initial “all-up” state [37, 38, 39, 40, 41, 42].
By tuning the pitch $q$ through the manipulation of the strength $B$ and/or the application time $\Delta t$ of the field gradient, different measurement setups can be prepared for each spiral plane. When denoting the number of different pitch angles by $n_{q}$, the number of independent experimental setups is $3n_{q}$, where the factor of 3 comes from the three spiral planes. For each setup, $2^{N}-1$ mutually independent expectation values can be obtained, as it is possible to choose whether or not to include the measurement of each individual qubit in the calculation of the multi-site correlation functions. In other words, from a single setup, say $\tilde{M}^{XY}(q)$ with a specific $q$, one can simultaneously obtain $\langle\tilde{M}^{XY}(q)\rangle$ and all the correlation functions where any of its $Z^{\prime}_{i}$s are replaced with $I$. We denote the corresponding operators by $\tilde{M}^{\alpha\beta}_{\bm{a}}(q)$, where $\alpha\beta=XY,YZ,ZX$ and $\bm{a}\in\{0,1\}^{N}$ is a multi-index indicating the positions of the identity matrices (labeled by 0), e.g., $\tilde{M}^{XY}_{(1,0,\ldots,0)}(q)=Z^{(1)\prime}\otimes I\cdots\otimes I/\sqrt{d}$, $\tilde{M}^{YZ}_{(1,1,0,\ldots,0)}(q)=Z^{(1)\prime\prime}\otimes Z^{(2)\prime\prime}\otimes I\cdots\otimes I/\sqrt{d}$, $\tilde{M}^{ZX}_{(1,\ldots,1)}(q)=Z^{(1)\prime\prime\prime}\otimes\cdots\otimes Z^{(N)\prime\prime\prime}/\sqrt{d}=\tilde{M}^{ZX}(q)$, etc. Note that this family includes the product of only identities at all sites, which contains no information. Consequently, the number of independent measurements that can be obtained from the three spiral sets with $n_{q}$ different pitch angles is, in total, $\sim 3n_{q}2^{N}$, which can be made comparable to $\mathcal{O}(rN^{2}2^{N})$, the required number of Pauli measurements for successful compressed sensing [20], by choosing a sufficiently large $n_{q}$.
@@
II.2 Numerical tests
To demonstrate the efficiency and accuracy of the spiral QST, we test the reconstruction of random pure states ($r=1$) and rank-3 mixed states ($r=3$). For the pitch angle $q$, we employ a set of $2N$ values with intervals of $\pi/N$, defined as
@@ II.2 Numerical tests
The above parameters, while reasonable from a physical standpoint, are by no means the only possible choice. We utilize the singular value thresholding (SVT) algorithm [20, 43] for compressed sensing, sampling expectation values of spiral operators in the set $\{\tilde{M}^{\alpha\beta}_{\bm{a}}(q)\}$ associated to $n_{q}$ randomly chosen pitch angles from Eq. (4) (see Appendix A). Target states $\rho$ are generated by randomly selecting from the Haar measure on a $d\times r$-dimensional system ($d=2^{N}$) and tracing out the $r$-dimensional ancilla. To model typical noise during the preparation of the target state, we add depolarizing noise of strength $\gamma$ after generating the target states as $\rho\rightarrow(1-\gamma)\rho+\gamma I^{\otimes N}/d$. We also account for measurement noise modeled by Gaussian fluctuations with standard deviation $\sigma$ in calculating expectation values for each measurement: $\langle\tilde{M}^{\alpha\beta}_{\bm{a}}(q)\rangle\rightarrow\mathcal{N}(\mu=\langle\tilde{M}^{\alpha\beta}_{\bm{a}}(q)\rangle,\sigma)$
Figure 2: Numerical tests of spiral QST. Sample-averaged fidelity and trace distance for compressed-sensing tomography of (a) random pure states (rank $r=1$) and (b) random mixed states (rank $r=3$) of 8 qubits, plotted as functions of the number of independent measurements $m$ scaled by $d^{2}=4^{8}$. The sampled states include 5 percent depolarizing noise ($\gamma=0.05$) and measurement Gaussian statistical noise with a standard deviation of $\sigma=0.1/d$. Spiral and Pauli measurements are compared in each case. For $r=3$, the results of Pauli-based tomography under the same conditions were also reported in previous work [20].
@@
[…]
where $A||_{1}\equiv{\rm Tr}(\sqrt{A^{\dagger}A})$ is the trace norm (or nuclear norm). These metrics take values 1 and 0, respectively, if and only if the two states $\rho_{1}$ and $\rho_{2}$ are identical. In Figs. 2(a) and 2(b), we present $F(\rho,\rho^{\rm rec})$ and $T(\rho,\rho^{\rm rec})$ between the target state $\rho$ and the reconstructed state $\rho^{\rm rec}$ for pure and mixed states, respectively, of $N=8$ qubits, plotted as functions of the number of spiral measurements (expectation values), $m$, used in the SVT algorithm, expressed in units of $d^{2}$. The results include noise levels corresponding to $\gamma=0.05$ and $\sigma=0.1/d$, matching those used in the demonstrations with Pauli-basis measurements in Ref. [20].
As seen from the results for $r=1$, the performance of spiral QST is comparable to (though slightly worse than) the conventional Pauli-basis compressed sensing [20]. The fidelity and trace distance indicate that the state reconstruction is very successful with only about 10 percent of a tomographically complete set of measurements. For mixed states with rank $r=3$, the results are also notable. Given that spiral measurements do not require precise single-site addressing nor a large number of experimental setups, this outcome highlights the practicality of our approach. Building on these results for random quantum states, we now aim to explore more realistic scenarios reflecting actual experimental conditions in the remainder of the paper.
Before that, let us note that one might be interested in comparing $\rho^{\rm rec}$ not with the ideal reference state $\rho$, but with the actual experimentally prepared state, which, in real situations, is affected by many sources of noise and therefore unknown. In this case one can make use of the “minimax” method for fidelity estimation developed in Refs. [44, 45]. Though it assumes a pure reference state, it does not impose any restrictions on the measurement operators (such as random drawing from a basis of Hermitian operators [46]), and it is possible to simply employ some of the spiral operators that were not used for the tomographic reconstruction; also, the sample complexity is comparable to that of the renowned “direct fidelity estimation” method [46]. Even if $\rho^{\rm rec}$ is mixed, we can still try to find a pure state that is close to it in fidelity to employ in the minimax method and then use the triangle inequality to estimate the fidelity between $\rho^{\rm rec}$ and the experimentally prepared state.
III Simulation of real-world experiments
III.1 Spiral-QST for typical ground states